At a certain college a random sample of 90 college students were surveyed and 75 students were found to have cars on campus. It had long been believed that 80% of students kept their cars on campus. Is there evidence, using alpha = 0.05, to say the proportion has increased?

  • State the type of test you are performing (multiple choice)
  • State the null and alternative hypotheses
  • Using Minitab results, state the test statistic
  • Using Minitab results, state the p-value
  • Decision
  • Conclusion
Answer #1

Ans:

z test for one proportion.

Ho:p=0.80

Ha . P 0.50

sample proportion=75/90=0.8333

Test statistic:

z=0.790

p-value=2*P(z>0.79)=0.4295

As,p-value>0.05,we fail to reject the null hypothesis.

There is not sufficient evidence to conclude that proportion of student kept their cars on campus is different from 0.8.